# Tree-Wasserstein Distance for High Dimensional Data with a Latent Feature Hierarchy

### v2

@@ D.1 Implementation Details
Setup. In our experiments, the datasets are divided into a $70\%$ training set and a $30\%$ testing set. The split of the word-document datasets in Sec. 5.2 follows previous work [Kusner et al., 2015, Huang et al., 2016]. This split was applied to both learning of the underlying feature tree and to the kNN model. Therefore, there is no train data leakage to the test data. The feature tree constructed from the training data can be reused to process new samples. For each evaluation, we employ a $k$NN classifier with varying $k\in\{1,3,5,7,9,11,13,15,17,19\}$. The evaluation is repeated five times, and we report the best result, averaged over these five runs.
@@
Datasets. In real-world word-document datasets, we test four word-document benchmarks in Kusner et al. [2015]: (i) the BBCSPORT consisting of 13243 bags of words (BOW) and 517 articles categorized into five sports types, (ii) the TWITTER comprising 6344 BOW and 2176 tweets in three types of sentiment, (iii) the CLASSIC, including 24277 BOW and 4965 academic papers from four publishers, and (iv) the AMAZON containing 42063 BOW and 5600 reviews of four products. The Word2Vec embedding [Mikolov et al., 2013] is used as the word embedding vectors that are trained on the Google News [^1] https://code.google.com/archive/p/word2vec/ dataset. The types of documents are used as labels in classification tasks. In scRNA-seq data analysis, we test two scRNA-seq datasets in Dumitrascu et al. [2021], where the datasets are available at the link [^2] https://github.com/solevillar/scGeneFit-python/tree/62f88ef0765b3883f592031ca593ec79679a52b4/scGeneFit/data_files. The first dataset, the Zeisel dataset, focuses on the mouse cortex and hippocampus [Zeisel et al., 2015]. It comprises 4000 gene markers and 3005 single cells. The second dataset, the CBMC dataset, is derived from a cord blood mononuclear cell study [Stoeckius et al., 2017]. It comprises 500 gene markers and 8617 single cells. We apply the divisive biclustering method from Zeisel et al. [2015] to obtain 47 subclasses for the Zeisel dataset and 56 subclasses for the CBMC dataset. The cell subclasses are used as labels in cell classification tasks. The Gene2Vec embeddings [Du et al., 2019] are used as the gene embedding vectors following the study of Gene Mover’s Distance [Bellazzi et al., 2021].
@@
Baselines. We compare the proposed TWD with existing TWDs, including (i) Quadtree [Indyk and Thaper, 2003] that constructs a random tree by a recursive division of hypercubes, (ii) Flowtree [Backurs et al., 2020] that computes the optimal flow of Quadtree, (iii) tree-sliced Wasserstein distance (TSWD) [Le et al., 2019] that computes an average of TWDs by random sampling trees, where the numbers of sampling are set to 1, 5, and 10, (iv) UltraTree [Chen et al., 2024] that constructs an ultrametric tree [Johnson, 1967] by minimizing OT regression cost, (v) weighted cluster TWD (WCTWD) [Yamada et al., 2022], (vi) weighted Quadtree TWD (WQTWD) [Yamada et al., 2022], and their sliced variants (vii) SWCTWD and (viii) SWQTWD [Yamada et al., 2022]. We also compare to TWDs where the trees are constructed by (i) MST [Prim, 1957], (ii) Tree Representation (TR) [Sonthalia and Gilbert, 2020] through a divide-and-conquer approach, (iii) gradient-based hierarchical clustering (HC) in hyperbolic space (gHHC) [Monath et al., 2019], (iv) gradient-based Ultrametric Fitting (UltraFit) [Chierchia and Perret, 2019], and (v) HC by hyperbolic Dasgupta’s cost (HHC) [Chami et al., 2020]. We refer to App. B for an extended description of these baselines. For non-TWD baselines, we in addition include the Word Mover’s Distance (WMD) [Kusner et al., 2015] for word-document data and we include the Gene Mover’s Distance (GMD) [Bellazzi et al., 2021] for scRNA-seq data. In the experiments in Sec. 5, the Euclidean distance $\widetilde{d}$ is used as the ground metric for WMD, GMD, and the TWD baselines, which aligns with existing studies [Kusner et al., 2015, Bellazzi et al., 2021, Indyk and Thaper, 2003, Backurs et al., 2020, Le et al., 2019, Chen et al., 2024, Yamada et al., 2022]. For the ablation study that uses cosine distance as the ground metric and directly performs classification tasks in the feature space using Euclidean distance and cosine distance, we refer to App. E.
@@
The maximal scale $K_{c}$ determines the range of scales over which the hyperbolic manifold operates. We follow [Lin et al., 2023] and suggest setting $K_{c}\in\{0,1,\ldots,19\}$. Our empirical study, consistent with [Lin et al., 2023], shows that increasing the maximal scale does not significantly impact classification accuracy. This robustness simplifies the parameter selection process by indicating that the method performs reliably across a broad range of $K_{c}$ values.
As these parameters are task-specific and do not have a one-size-fits-all solution, we use Optuna [Akiba et al., 2019] to efficiently explore the parameter space and identify optimal settings, where $(\epsilon,K_{c})$ are set $(10,12)$, $(5,7)$, $(1,15)$, $(10,14)$, $(2,7)$, and $(5,9)$ for BBCSPORT, TWITTER, CLASSIC, AMAZON, Zeisel, and CBMC, respectively. We note that the selection of kernel type, scale, and the initial distance metric in the kernel influences the effectiveness of our approach. Such choices are not only crucial but also task-specific, which is in line with the common practice in kernel and manifold learning methods. These elements, which do not have a one-size-fits-all solution, continue to be of wide interest in ongoing research [Lindenbaum et al., 2020a].
In our study, we employed the Gaussian kernel with a distance measure based on cosine similarity following prior research [Jaskowiak et al., 2014, Kenter and De Rijke, 2015], which is standard and commonly used in classification tasks and manifold learning. In our empirical analysis, we observed that employing the Euclidean distance in the ambient space tends to be highly sensitive and underperforms compared to using a distance based on the cosine similarity in the tested datasets, as detailed in the App. E. Our primary objective is to highlight the unique aspects of our method rather than kernel selection, which, while crucial, is a common consideration across all kernel and manifold learning methods.
@@ D.3 Accommodating Large Datasets
In our work, we note that constructing the diffusion kernels and their eigenvectors by applying the eigendecomposition to them is the most computationally intensive step. However, recent methods in diffusion geometry have proposed various techniques to significantly reduce the run time and space complexity of diffusion operators. Here, we consider the diffusion landmark [Shen and Wu, 2022] for scalability improvements which reduces the complexity to $O(n^{1+2\tau})$ from the much larger $O(n^{3})$, where $\tau<1$ is the proportion of the landmark set. We briefly describe the diffusion landmark approach below.
Consider a set of points $\mathcal{A}=\{\mathbf{a}_{j}\}_{j=1}^{m}\subseteq\mathbb{R}^{n}$, let $\mathcal{A}^{\prime}=\{\mathbf{a}_{j}^{\prime}\}_{j=1}^{m^{\prime}}\subseteq\mathbb{R}^{n}$ be a subset of $\mathcal{A}$, where $1>\tau=\log_{m}(m^{\prime})$. A landmark-set affinity matrix is constructed by $\mathbf{K}(j,j^{\prime})=\exp(-d^{2}(j,j^{\prime})/\epsilon)$, where $d(j,j^{\prime})$ denotes a suitable distance between the data points $\mathbf{a}_{j}\in\mathcal{A}$ and $\mathbf{a}_{j}^{\prime}\in\mathcal{A}^{\prime}$. Let $\widetilde{\mathbf{D}}$ be a diagonal matrix, where the diagonal elements are $\widetilde{\mathbf{D}}(j,j)=\mathbf{e}_{j}^{\top}\mathbf{KK}^{\top}\mathbf{1}_{m}$. The eigen-structure of an SPD matrix $\mathbf{D}^{-1/2}\widetilde{\mathbf{Q}}\mathbf{D}^{-1/2}$ that is similar to the diffusion operator [Coifman and Hirn, 2014, Katz et al., 2020] is recovered by applying SVD to the matrix $\widetilde{\mathbf{D}}^{-1/2}\mathbf{K}=\widetilde{\mathbf{U}}\widetilde{\mathbf{\Lambda}}\widetilde{\mathbf{V}}$, where $\widetilde{\mathbf{Q}}=\mathbf{KK}^{\top}\in\mathbb{R}^{m\times m}$ is the landmark-affinity matrix. The technique of constructing the diffusion on $\mathcal{A}$ by going through the subset $\mathcal{A}^{\prime}$ and their eigenvectors can be seamlessly integrated into our TWD, enabling the analysis of datasets larger than ten thousand data points, e.g., the word-document data in Sec. 5.2.
@@ E.1 Toy Problem
Consider a four-level balanced binary tree $T=(V,E,\mathbf{A})$ rooted at node $v_{1}$, where $V=\{v_{1},\ldots,v_{15}\}$ is the vertex set organized from the root to the leaves of the tree, $E$ is the edge set, and $\mathbf{A}\in\mathbb{R}^{15\times 15}$ is the non-negative edge weights. Consider a set of vectors $\{\mathbf{x}_{i}\in\mathbb{R}^{15}\}_{i=1}^{n}$ such that $\mathbf{x}_{i}$ is a non-negative realization from a normal distribution $\mathcal{N}(\bm{\mu},\mathbf{L}^{\dagger})$ [Rue and Held, 2005], where $\bm{\mu}=5\cdot\mathbf{1}_{15}\in\mathbb{R}^{15}$ and $\mathbf{L}$ is the graph Laplacian of $T$ [^3] The mean is large compared to the covariance matrix, and the realizations are non-negative with high probability. Realizations with negative elements are rejected..
@@ E.2 Runtime Analysis
In Fig. 4, we present the runtime performance of our TWD in comparison to other TWDs and Wasserstein distances when applied to scRNA-seq and word-document datasets. Although our TWD is not the fastest method, it consistently ranks as the fourth or fifth quickest. It outperforms deep-based methods like HHC-TWD and gHHC-TWD, as well as pre-trained OT baselines such as GMD for single-cell RNA sequencing datasets and WMD for word-document datasets. The fastest methods identified are TSWD-1, TR-TWD, and MST-TWD. The work in [Le et al., 2019] reports that the runtime of TSWD is influenced by the number of tree slices used in sampling. Similarly, slice variants SWCTWD and SWQTWD are also affected by the number of slices. Despite being slightly slower compared to TSWD-1, TR-TWD, and MST-TWD, our TWD demonstrates significant advantages in terms of classification accuracy, as shown in Tab. LABEL:tab:classification. We also report the runtime of the OT distance using the embedding distance between features as a ground distance, denoted as OT-ED. We note that this computational expense is $O(mn^{3}+m^{3}\log m)$. In contrast, our TWD offers a substantially reduced runtime of $O(m^{1.2})$, thereby demonstrating a significant advantage when handling large datasets. Note that the scalability of the proposed TWD is enhanced by utilizing diffusion landmarks [Shen and Wu, 2022], as detailed in App. D.3. This enhancement allows our TWD to achieve computational performance comparable to TSWD-1, TR-TWD, and MST-TWD, particularly in datasets with a large number of features, such as the word-document datasets shown in Fig. 4.
@@ E.3.2 TWD by MST and AT
Fig. 6 shows the classification performance of the scRNA-seq and word-document datasets using the TWD with the tree constructed using Minimum Spanning Trees (MST) [Prim, 1957] and Approximating Tree (AT) [Chepoi et al., 2008]. To ensure that the tree nodes are contained in a subtree, which is required when computing the TWD in Eq. (1), we set the minimal leaf number as the number of features in MST for computing the TWD between samples. We see that the TWD approximated by either the MST or the AT is less effective than our TWD, implying that MST and AT do not capture well the underlying tree structure between features compared to our binary tree construction.
@@ E.3.3 Tree Construction by Embedding Distance
We note that the embedding distance [Lin et al., 2023] could be simply used for tree construction in existing TWDs [Indyk and Thaper, 2003, Le et al., 2019, Yamada et al., 2022], graph-based algorithms [Alon et al., 1995], or linkage methods [Murtagh and Contreras, 2012] instead of the HD-LCA in the tree decoding (Alg. 1). However, these generic algorithms, while providing upper bounds on the distortion of the constructed tree distances, do not build a geometrically meaningful tree when given a tree metric [Sonthalia and Gilbert, 2020], because they do not exploit the hyperbolic geometry of the embedded space or do not have theoretical guarantees to recover the hidden feature hierarchy.
@@ E.3.4 Exploring Other Hyperbolic Embedding and the Impact of Curvature
Hyperbolic embedding methods [Chamberlain et al., 2017, Nickel and Kiela, 2017, 2018] typically assume a fully known or partially known graph of $T$, which in turn is embedded into hyperbolic space. In Sonthalia and Gilbert [2020], Lin et al. [2023], Sala et al. [2018], Weber et al. [2024], these methods were shown to inadequately capture the intrinsic hierarchical relation underlying high-dimensional observational data $\mathbf{X}$ (i.e., the node attributes). Therefore, we opted to use hyperbolic embedding [Lin et al., 2023], a recent hyperbolic embedding method with theoretical guarantees. This embedding and its accompanying theory enabled us to introduce our theoretical results in Sec. 4.
@@
It is important to note that we could indeed apply traditional hyperbolic embedding methods directly to the affinity matrix $\mathbf{Q}$. From a graph perspective, this is the affinity of a graph, but this graph is not usually a tree. These traditional methods, due to the quantity they are minimizing, will estimate a hyperbolic distance that approximates as best as possible the shortest path distance on this non-tree graph induced by $\mathbf{Q}$. As such, they do not, in general, recover a hyperbolic distance related to the hyperbolic distance of the hidden tree $T$. In contrast, the hyperbolic diffusion embedding approach [Lin et al., 2023] is designed such that the recovered hyperbolic metric is close to that of the hidden tree $T$, guaranteeing it will be bilipschitz equivalent to it. One of the reasons for the success of this method is that the obtained embedding is not provided by the minimization of an objective function controlled by $\mathbf{Q}$. Instead, Lin [Lin et al., 2023] designed a non-trivial process that is neither iterative nor based on optimization that computes the embedding directly with theoretical guarantees. We build on the theoretical guarantees of this embedding in our own theoretical contribution in Thm. 4.1.
@@
We here give an intuition as to the reason for the success of the approach from Lin et al. [2023]. After normalizing the affinity matrix $\mathbf{Q}$, the diffusion operator $\mathbf{P}$ is computed. This operator was proven to reveal the underlying manifold. This property is what allows an embedding approach based on the diffusion operator to approximately recover the underlying tree metric $d_{T}$ rather than fit the metric of a non-tree graph constructed from the affinity matrix $\mathbf{Q}$. The diffusion operator reveals the underlying manifold in the following sense [Coifman and Lafon, 2006]. In the limit of the infinite number of features $m\rightarrow\infty$ and small scale $\epsilon\rightarrow 0$, the operator $\mathbf{P}^{t/\epsilon}$ pointwise converges to the Neumann heat kernel of the underlying manifold $\mathbf{H}_{t}=\exp(-t\Delta)$, where $\Delta$ is the Laplace–Beltrami operator on the manifold. In other words, the diffusion operator is a discrete approximation of the heat kernel on the manifold. Intuitively, $\mathbf{P}^{t}$ is designed to reveal the local connectivity at diffusion time scale $t$.
@@
Considering multiple timescales on a dyadic grid $\{2^{-k}\}_{k=0}^{K_{c}}$ allows for the association of different diffusion timescales with neighborhoods of varying sizes. The multi-scale embedding of the associated diffusion operators in hyperbolic space generates distances that are exponentially scaled, aligning with the structure of a tree distance. This scaling naturally introduces a hierarchical relationship between the diffusion timescales, enabling the estimation of a latent hierarchical metric. From a theoretical standpoint, it was shown that the $\ell_{1}$ distance in this multi-scale embedding induced by the metric of the product of hyperbolic spaces can estimate the distance $d_{T}$ in the hidden hierarchical space that underlies the observations even when $d_{T}$ is not given or when we do not have access to the explicit $T$ [Lin et al., 2023]. It is stated under the assumption that $(\mathbf{P}^{t})_{t\in(0,1]}$ is a pointwise approximation of the heat kernel. Such an approximation, mainly in the limit $m\rightarrow\infty$ and $\epsilon\rightarrow 0$, was shown and studied in Coifman and Lafon [2006], Singer [2006], Belkin and Niyogi [2008]. In addition, three strong regularity conditions are required: an upper bound on the operator elements, a lower bound on the operator elements, and a Hölder continuity condition. Importantly, the heat kernel satisfies these conditions [Lin et al., 2023].
@@
In our work, we use the hyperbolic diffusion embedding from Lin et al. [2023] to present a novel data-driven tree decoding method that operates in high-dimensional hyperbolic spaces. Our method is based on a new definition of the counterpart of the LCA in hyperbolic spaces. Our Thm. 4.1 is built on the multi-scale hyperbolic embedding from Lin et al. [2023], where we use the theoretical property [Lin et al., 2023] that the $\ell_{1}$ distance in this embedding can estimate the hidden tree metric $d_{T}$. The difference between Thm. 4.1 and the work from Lin et al. [2023] is that we construct an explicit tree $B$ whose tree metric approximates the hidden tree metric $d_{T}$. In Lin et al. [2023], no tree is recovered, and only the hidden tree metric is approximated by the hyperbolic embedding distances. Our Thm. 4.2 is established based on Thm. 4.1, where the proposed TWD can estimate the TWD associated with the latent feature tree $T$ without access to $T$. The uniqueness of our approach lies in its ability to work without the need for a pre-known hierarchical graph.

### v3

@@ D.1 Implementation Details
Setup. In our experiments, the datasets are divided into a $70\%$ training set and a $30\%$ testing set. The split of the word-document datasets in Sec. 5.2 follows previous work (Kusner et al., 2015; Huang et al., 2016). This split was applied to both learning of the underlying feature tree and to the kNN model. Therefore, there is no train data leakage to the test data. The feature tree constructed from the training data can be reused to process new samples. For each evaluation, we employ a $k$NN classifier with varying $k\in\{1,3,5,7,9,11,13,15,17,19\}$. The evaluation is repeated five times, and we report the best result, averaged over these five runs.
@@
Datasets. In real-world word-document datasets, we test four word-document benchmarks in Kusner et al. (2015): (i) the BBCSPORT consisting of 13243 bags of words (BOW) and 517 articles categorized into five sports types, (ii) the TWITTER comprising 6344 BOW and 2176 tweets in three types of sentiment, (iii) the CLASSIC, including 24277 BOW and 4965 academic papers from four publishers, and (iv) the AMAZON containing 42063 BOW and 5600 reviews of four products. The Word2Vec embedding (Mikolov et al., 2013) is used as the word embedding vectors that are trained on the Google News [^1] https://code.google.com/archive/p/word2vec/ dataset. The types of documents are used as labels in classification tasks. In scRNA-seq data analysis, we test two scRNA-seq datasets in Dumitrascu et al. (2021), where the datasets are available at the link [^2] https://github.com/solevillar/scGeneFit-python/tree/62f88ef0765b3883f592031ca593ec79679a52b4/scGeneFit/data_files. The first dataset, the Zeisel dataset, focuses on the mouse cortex and hippocampus (Zeisel et al., 2015). It comprises 4000 gene markers and 3005 single cells. The second dataset, the CBMC dataset, is derived from a cord blood mononuclear cell study (Stoeckius et al., 2017). It comprises 500 gene markers and 8617 single cells. We apply the divisive biclustering method from Zeisel et al. (2015) to obtain 47 subclasses for the Zeisel dataset and 56 subclasses for the CBMC dataset. The cell subclasses are used as labels in cell classification tasks. The Gene2Vec embeddings (Du et al., 2019) are used as the gene embedding vectors following the study of Gene Mover’s Distance (Bellazzi et al., 2021).
@@
Baselines. We compare the proposed TWD with existing TWDs, including (i) Quadtree (Indyk & Thaper, 2003) that constructs a random tree by a recursive division of hypercubes, (ii) Flowtree (Backurs et al., 2020) that computes the optimal flow of Quadtree, (iii) tree-sliced Wasserstein distance (TSWD) (Le et al., 2019) that computes an average of TWDs by random sampling trees, where the numbers of sampling are set to 1, 5, and 10, (iv) UltraTree (Chen et al., 2024) that constructs an ultrametric tree (Johnson, 1967) by minimizing OT regression cost, (v) weighted cluster TWD (WCTWD) (Yamada et al., 2022), (vi) weighted Quadtree TWD (WQTWD) (Yamada et al., 2022), and their sliced variants (vii) SWCTWD and (viii) SWQTWD (Yamada et al., 2022). We also compare to TWDs where the trees are constructed by (i) MST (Prim, 1957), (ii) Tree Representation (TR) (Sonthalia & Gilbert, 2020) through a divide-and-conquer approach, (iii) gradient-based hierarchical clustering (HC) in hyperbolic space (gHHC) (Monath et al., 2019), (iv) gradient-based Ultrametric Fitting (UltraFit) (Chierchia & Perret, 2019), and (v) HC by hyperbolic Dasgupta’s cost (HHC) (Chami et al., 2020). We refer to App. B for an extended description of these baselines. For non-TWD baselines, we in addition include the Word Mover’s Distance (WMD) (Kusner et al., 2015) for word-document data and we include the Gene Mover’s Distance (GMD) (Bellazzi et al., 2021) for scRNA-seq data. In the experiments in Sec. 5, the Euclidean distance $\widetilde{d}$ is used as the ground metric for WMD, GMD, and the TWD baselines, which aligns with existing studies (Kusner et al., 2015; Bellazzi et al., 2021; Indyk & Thaper, 2003; Backurs et al., 2020; Le et al., 2019; Chen et al., 2024; Yamada et al., 2022). For the ablation study that uses cosine distance as the ground metric and directly performs classification tasks in the feature space using Euclidean distance and cosine distance, we refer to App. E.
@@
The maximal scale $K_{c}$ determines the range of scales over which the hyperbolic manifold operates. We follow (Lin et al., 2023) and suggest setting $K_{c}\in\{0,1,\ldots,19\}$. Our empirical study, consistent with (Lin et al., 2023), shows that increasing the maximal scale does not significantly impact classification accuracy. This robustness simplifies the parameter selection process by indicating that the method performs reliably across a broad range of $K_{c}$ values.
As these parameters are task-specific and do not have a one-size-fits-all solution, we use Optuna (Akiba et al., 2019) to efficiently explore the parameter space and identify optimal settings, where $(\epsilon,K_{c})$ are set $(10,12)$, $(5,7)$, $(1,15)$, $(10,14)$, $(2,7)$, and $(5,9)$ for BBCSPORT, TWITTER, CLASSIC, AMAZON, Zeisel, and CBMC, respectively. We note that the selection of kernel type, scale, and the initial distance metric in the kernel influences the effectiveness of our approach. Such choices are not only crucial but also task-specific, which is in line with the common practice in kernel and manifold learning methods. These elements, which do not have a one-size-fits-all solution, continue to be of wide interest in ongoing research (Lindenbaum et al., 2020a).
In our study, we employed the Gaussian kernel with a distance measure based on cosine similarity following prior research (Jaskowiak et al., 2014; Kenter & De Rijke, 2015), which is standard and commonly used in classification tasks and manifold learning. In our empirical analysis, we observed that employing the Euclidean distance in the ambient space tends to be highly sensitive and underperforms compared to using a distance based on the cosine similarity in the tested datasets, as detailed in the App. E. Our primary objective is to highlight the unique aspects of our method rather than kernel selection, which, while crucial, is a common consideration across all kernel and manifold learning methods.
@@ D.3 Accommodating Large Datasets
In our work, we note that constructing the diffusion kernels and their eigenvectors by applying the eigendecomposition to them is the most computationally intensive step. However, recent methods in diffusion geometry have proposed various techniques to significantly reduce the run time and space complexity of diffusion operators. Here, we consider the diffusion landmark (Shen & Wu, 2022) for scalability improvements which reduces the complexity to $O(n^{1+2\tau})$ from the much larger $O(n^{3})$, where $\tau<1$ is the proportion of the landmark set. We briefly describe the diffusion landmark approach below.
Consider a set of points $\mathcal{A}=\{\mathbf{a}_{j}\}_{j=1}^{m}\subseteq\mathbb{R}^{n}$, let $\mathcal{A}^{\prime}=\{\mathbf{a}_{j}^{\prime}\}_{j=1}^{m^{\prime}}\subseteq\mathbb{R}^{n}$ be a subset of $\mathcal{A}$, where $1>\tau=\log_{m}(m^{\prime})$. A landmark-set affinity matrix is constructed by $\mathbf{K}(j,j^{\prime})=\exp(-d^{2}(j,j^{\prime})/\epsilon)$, where $d(j,j^{\prime})$ denotes a suitable distance between the data points $\mathbf{a}_{j}\in\mathcal{A}$ and $\mathbf{a}_{j}^{\prime}\in\mathcal{A}^{\prime}$. Let $\widetilde{\mathbf{D}}$ be a diagonal matrix, where the diagonal elements are $\widetilde{\mathbf{D}}(j,j)=\mathbf{e}_{j}^{\top}\mathbf{KK}^{\top}\mathbf{1}_{m}$. The eigen-structure of an SPD matrix $\mathbf{D}^{-1/2}\widetilde{\mathbf{Q}}\mathbf{D}^{-1/2}$ that is similar to the diffusion operator (Coifman & Hirn, 2014; Katz et al., 2020) is recovered by applying SVD to the matrix $\widetilde{\mathbf{D}}^{-1/2}\mathbf{K}=\widetilde{\mathbf{U}}\widetilde{\mathbf{\Lambda}}\widetilde{\mathbf{V}}$, where $\widetilde{\mathbf{Q}}=\mathbf{KK}^{\top}\in\mathbb{R}^{m\times m}$ is the landmark-affinity matrix. The technique of constructing the diffusion on $\mathcal{A}$ by going through the subset $\mathcal{A}^{\prime}$ and their eigenvectors can be seamlessly integrated into our TWD, enabling the analysis of datasets larger than ten thousand data points, e.g., the word-document data in Sec. 5.2.
@@ E.1 Toy Problem
Consider a four-level balanced binary tree $T=(V,E,\mathbf{A})$ rooted at node $v_{1}$, where $V=\{v_{1},\ldots,v_{15}\}$ is the vertex set organized from the root to the leaves of the tree, $E$ is the edge set, and $\mathbf{A}\in\mathbb{R}^{15\times 15}$ is the non-negative edge weights. Consider a set of vectors $\{\mathbf{x}_{i}\in\mathbb{R}^{15}\}_{i=1}^{n}$ such that $\mathbf{x}_{i}$ is a non-negative realization from a normal distribution $\mathcal{N}(\bm{\mu},\mathbf{L}^{\dagger})$ (Rue & Held, 2005), where $\bm{\mu}=5\cdot\mathbf{1}_{15}\in\mathbb{R}^{15}$ and $\mathbf{L}$ is the graph Laplacian of $T$ [^3] The mean is large compared to the covariance matrix, and the realizations are non-negative with high probability. Realizations with negative elements are rejected..
@@ E.2 Runtime Analysis
In Fig. 4, we present the runtime performance of our TWD in comparison to other TWDs and Wasserstein distances when applied to scRNA-seq and word-document datasets. Although our TWD is not the fastest method, it consistently ranks as the fourth or fifth quickest. It outperforms deep-based methods like HHC-TWD and gHHC-TWD, as well as pre-trained OT baselines such as GMD for single-cell RNA sequencing datasets and WMD for word-document datasets. The fastest methods identified are TSWD-1, TR-TWD, and MST-TWD. The work in (Le et al., 2019) reports that the runtime of TSWD is influenced by the number of tree slices used in sampling. Similarly, slice variants SWCTWD and SWQTWD are also affected by the number of slices. Despite being slightly slower compared to TSWD-1, TR-TWD, and MST-TWD, our TWD demonstrates significant advantages in terms of classification accuracy, as shown in Tab. LABEL:tab:classification. We also report the runtime of the OT distance using the embedding distance between features as a ground distance, denoted as OT-ED. We note that this computational expense is $O(mn^{3}+m^{3}\log m)$. In contrast, our TWD offers a substantially reduced runtime of $O(m^{1.2})$, thereby demonstrating a significant advantage when handling large datasets. Note that the scalability of the proposed TWD is enhanced by utilizing diffusion landmarks (Shen & Wu, 2022), as detailed in App. D.3. This enhancement allows our TWD to achieve computational performance comparable to TSWD-1, TR-TWD, and MST-TWD, particularly in datasets with a large number of features, such as the word-document datasets shown in Fig. 4.
@@ E.3.2 TWD by MST and AT
Fig. 6 shows the classification performance of the scRNA-seq and word-document datasets using the TWD with the tree constructed using Minimum Spanning Trees (MST) (Prim, 1957) and Approximating Tree (AT) (Chepoi et al., 2008). To ensure that the tree nodes are contained in a subtree, which is required when computing the TWD in Eq. (1), we set the minimal leaf number as the number of features in MST for computing the TWD between samples. We see that the TWD approximated by either the MST or the AT is less effective than our TWD, implying that MST and AT do not capture well the underlying tree structure between features compared to our binary tree construction.
@@ E.3.3 Tree Construction by Embedding Distance
We note that the embedding distance (Lin et al., 2023) could be simply used for tree construction in existing TWDs (Indyk & Thaper, 2003; Le et al., 2019; Yamada et al., 2022), graph-based algorithms (Alon et al., 1995), or linkage methods (Murtagh & Contreras, 2012) instead of the HD-LCA in the tree decoding (Alg. 1). However, these generic algorithms, while providing upper bounds on the distortion of the constructed tree distances, do not build a geometrically meaningful tree when given a tree metric (Sonthalia & Gilbert, 2020), because they do not exploit the hyperbolic geometry of the embedded space or do not have theoretical guarantees to recover the hidden feature hierarchy.
@@ E.3.4 Exploring Other Hyperbolic Embedding and the Impact of Curvature
Hyperbolic embedding methods (Chamberlain et al., 2017; Nickel & Kiela, 2017; 2018) typically assume a fully known or partially known graph of $T$, which in turn is embedded into hyperbolic space. In Sonthalia & Gilbert (2020); Lin et al. (2023); Sala et al. (2018); Weber et al. (2024), these methods were shown to inadequately capture the intrinsic hierarchical relation underlying high-dimensional observational data $\mathbf{X}$ (i.e., the node attributes). Therefore, we opted to use hyperbolic embedding (Lin et al., 2023), a recent hyperbolic embedding method with theoretical guarantees. This embedding and its accompanying theory enabled us to introduce our theoretical results in Sec. 4.
@@
It is important to note that we could indeed apply traditional hyperbolic embedding methods directly to the affinity matrix $\mathbf{Q}$. From a graph perspective, this is the affinity of a graph, but this graph is not usually a tree. These traditional methods, due to the quantity they are minimizing, will estimate a hyperbolic distance that approximates as best as possible the shortest path distance on this non-tree graph induced by $\mathbf{Q}$. As such, they do not, in general, recover a hyperbolic distance related to the hyperbolic distance of the hidden tree $T$. In contrast, the hyperbolic diffusion embedding approach (Lin et al., 2023) is designed such that the recovered hyperbolic metric is close to that of the hidden tree $T$, guaranteeing it will be bilipschitz equivalent to it. One of the reasons for the success of this method is that the obtained embedding is not provided by the minimization of an objective function controlled by $\mathbf{Q}$. Instead, Lin (Lin et al., 2023) designed a non-trivial process that is neither iterative nor based on optimization that computes the embedding directly with theoretical guarantees. We build on the theoretical guarantees of this embedding in our own theoretical contribution in Thm. 4.1.
@@
We here give an intuition as to the reason for the success of the approach from Lin et al. (2023). After normalizing the affinity matrix $\mathbf{Q}$, the diffusion operator $\mathbf{P}$ is computed. This operator was proven to reveal the underlying manifold. This property is what allows an embedding approach based on the diffusion operator to approximately recover the underlying tree metric $d_{T}$ rather than fit the metric of a non-tree graph constructed from the affinity matrix $\mathbf{Q}$. The diffusion operator reveals the underlying manifold in the following sense (Coifman & Lafon, 2006). In the limit of the infinite number of features $m\rightarrow\infty$ and small scale $\epsilon\rightarrow 0$, the operator $\mathbf{P}^{t/\epsilon}$ pointwise converges to the Neumann heat kernel of the underlying manifold $\mathbf{H}_{t}=\exp(-t\Delta)$, where $\Delta$ is the Laplace–Beltrami operator on the manifold. In other words, the diffusion operator is a discrete approximation of the heat kernel on the manifold. Intuitively, $\mathbf{P}^{t}$ is designed to reveal the local connectivity at diffusion time scale $t$.
@@
Considering multiple timescales on a dyadic grid $\{2^{-k}\}_{k=0}^{K_{c}}$ allows for the association of different diffusion timescales with neighborhoods of varying sizes. The multi-scale embedding of the associated diffusion operators in hyperbolic space generates distances that are exponentially scaled, aligning with the structure of a tree distance. This scaling naturally introduces a hierarchical relationship between the diffusion timescales, enabling the estimation of a latent hierarchical metric. From a theoretical standpoint, it was shown that the $\ell_{1}$ distance in this multi-scale embedding induced by the metric of the product of hyperbolic spaces can estimate the distance $d_{T}$ in the hidden hierarchical space that underlies the observations even when $d_{T}$ is not given or when we do not have access to the explicit $T$ (Lin et al., 2023). It is stated under the assumption that $(\mathbf{P}^{t})_{t\in(0,1]}$ is a pointwise approximation of the heat kernel. Such an approximation, mainly in the limit $m\rightarrow\infty$ and $\epsilon\rightarrow 0$, was shown and studied in Coifman & Lafon (2006); Singer (2006); Belkin & Niyogi (2008). In addition, three strong regularity conditions are required: an upper bound on the operator elements, a lower bound on the operator elements, and a Hölder continuity condition. Importantly, the heat kernel satisfies these conditions (Lin et al., 2023).
@@
In our work, we use the hyperbolic diffusion embedding from Lin et al. (2023) to present a novel data-driven tree decoding method that operates in high-dimensional hyperbolic spaces. Our method is based on a new definition of the counterpart of the LCA in hyperbolic spaces. Our Thm. 4.1 is built on the multi-scale hyperbolic embedding from Lin et al. (2023), where we use the theoretical property (Lin et al., 2023) that the $\ell_{1}$ distance in this embedding can estimate the hidden tree metric $d_{T}$. The difference between Thm. 4.1 and the work from Lin et al. (2023) is that we construct an explicit tree $B$ whose tree metric approximates the hidden tree metric $d_{T}$. In Lin et al. (2023), no tree is recovered, and only the hidden tree metric is approximated by the hyperbolic embedding distances. Our Thm. 4.2 is established based on Thm. 4.1, where the proposed TWD can estimate the TWD associated with the latent feature tree $T$ without access to $T$. The uniqueness of our approach lies in its ability to work without the need for a pre-known hierarchical graph.